OpenStudy (pokemon23):

Find three consecutive integers such that the sum of the first two integers is 24 more than the third integers. Find three consecutive integers such that the sum of the first two integers is 24 more than the third integers. @Mathematics

OpenStudy (pokemon23):

Thanks Mert ;)

OpenStudy (mertsj):

the integers are x x+1 and x+2 The sum of the first two is 2x+1. That sum is bigger then x+2. In fact, it is 24 bigger. If you want to write that they are equal, you must either subtract 24 from the big one which is 2x+1 or add 24 to the little one which is x+2 2x+1-24 = x+2 2x-23 = x+2 x-23=2 x=25 x+1=27 x+2=28

OpenStudy (anonymous):

25, 26 & 27